Using the "Example Dataset" and SPSS, apply the t-test to assess the following statement: "Men and women have different incomes in this city."

  1. Describe what t-test is the most appropriate and explain why. Discuss whether you used a one-tailed or two-tailed test and explain why.
  2. Using SPSS, calculate the t-test and provide the test statistic and critical value assuming an alpha of .05.
  3. Calculate the effect size using r2.
Sex
Female
Male
Female
Male
Female
Female
Female
Male
Female
Male
Female
Male
Female
Male
Male
Male
Female
Male
Male
Female
Female
Female
Male
Male
Female
Male
Female
Female
Male
Male
Annual_Income*
51000
23000
35000
10000
28000
5000
46000
36000
51000
12000
78000
34000
15000
28000
28000
24000
55000
62000
32000
7000
17000
64000
5000
14000
20000
72000
85000
15000
64000
27000
Answer #1

A t-test for two means with unknown population variances and two independent samples is a hypothesis test that attempts to make a claim about the population means (μ2​).

More specifically, a t-test uses sample information to assess how plausible it is for the population means μ1​ and μ2​ to be equal. The test has two non-overlapping hypotheses, the null and the alternative hypothesis.

The formula for a t-statistic for two population means

Steps to be followed in SPSS

1) Enter the data in Data view with variable named as Annual Income and Gender

2) Go to analyse--> compare means-->independent sample t- test

3) Test variable = Annual_income, Grouping variable = Gender

4) define group as 1 and 2 for male and female.

5) click on ok

Annual_income

Gender

51000.00

Female

23000.00

Male

35000.00

Female

10000.00

Male

28000.00

Female

5000.00

Female

46000.00

Female

36000.00

Male

51000.00

Female

12000.00

Male

78000.00

Female

34000.00

Male

15000.00

Female

28000.00

Male

28000.00

Male

24000.00

Male

55000.00

Female

62000.00

Male

32000.00

Male

7000.00

Female

17000.00

Female

64000.00

Female

5000.00

Male

14000.00

Male

20000.00

Female

72000.00

Male

85000.00

Female

15000.00

Female

64000.00

Male

27000.00

Male

Group Statistics

Gender

N

Mean

Std. Deviation

Std. Error Mean

Annual_income

Male

15

31400.0000

20138.09468

5199.63369

Female

15

38133.3333

25575.84426

6603.65459

Independent Samples Test

Levene's Test for Equality of Variances

t-test for Equality of Means

F

Sig.

t

df

Sig. (2-tailed)

Mean Difference

Std. Error Difference

95% Confidence Interval of the Difference

Lower

Upper

Annual_income

Equal variances assumed

2.236

.146

-.801

28

.430

-6733.33333

8405.02495

-23950.24647

10483.57981

Equal variances not assumed

-.801

26.540

.430

-6733.33333

8405.02495

-23993.03136

10526.36470

             

From the above table p value for t-test for Equality of Means is 0.430 which is greater than 0.05 (i.e. 5% level of significance ) so we fail to reject the null hypothesis and conclude that there is no significant difference between the annual income of female and male.

Effective size

Cohen's d = (M2 - M1) ⁄ SDpooled

SDpooled = √((SD12 + SD22) ⁄ 2)

Cohen's d = (38133.33 - 31400) ⁄ 23018.104402 = 0.292523.

Gates' delta = (38133.33 - 31400) ⁄ 20138.09 = 0.334358.

Hedges' g = (38133.33 - 31400) ⁄ 23018.104402 = 0.292523.

answer